### v2

@@ E.2 Evaluation Datasets
For the evaluation on the TSLib benchmark, the datasets were split into training, validation, and test sets. The split for the ETT and Weather datasets follows the configuration adopted by iTransformer (Liu et al., 2023). All other datasets use a 70%/10%/20% ratio for the training, validation, and test sets, respectively.
E.3 Evaluation Length Selection
In this section, we provide further details regarding the selection of historical sequence lengths for the various models evaluated in the TSLib benchmark, supplementing the discussion in Section 4.1.1.
To accommodate diverse application scenarios, an increasing number of TSFMs (Liu et al., 2025; Das et al., 2024; Liu et al., 2024c; Woo et al., 2024) have devoted attention to predicting over long contexts. Consequently, we evaluate Kairos and other TSFMs under a long‑context setting. Specifically, we adopt a context length of 2048 time steps and examine four prediction horizons, which are {96,192,336,720}. For TSFMs incapable of processing this context length, we instead employ the context length at which each model achieves its best performance.
For the full-shot deep learning baselines, a hyperparameter search was conducted independently for each dataset to ensure a fair evaluation. We compared the performance of each model using the context length from its original publication against a length of 2048, selecting the superior of the two. To be specific, lengths of 96 and 2048 were compared for DLinear (Zeng et al., 2022), iTransformer (Liu et al., 2023), TimesNet (Wu et al., 2022), and Pathformer, while lengths of 336, 512, and 2048 were compared for PatchTST (Nie et al., 2023).
@@ E.3 Evaluation Length Selection
E.4 Details of DRoPE Analysis
In this section, we explain in more detail the setting of the DRoPE analysis experiment discussed in Section 4.4.2. To more definitively ascertain whether these instance-specific modulations truly capture and leverage beneficial positional information derived from an instance’s FFT features, we designed a series of ?shuffle? experiments. These experiments function as a form of causal intervention analysis. The underlying hypothesis is: if the instance-derived $\theta_{\text{inst}}$ modulations are crucial for DRoPE’s performance, then disrupting this linkage by applying $\theta_{\text{inst}}$ modulations from one instance to another (shuffling) should lead to a noticeable degradation in forecasting accuracy.
@@ E.4 Details of DRoPE Analysis
  • Fixed RoPE: We use the fully trained DRoPE model but disable its instance-adaptive modulation during inference. Specifically, the prediction of $\gamma^{(l)}$ and $\beta^{(l)}$ is bypassed, and all instances revert to using the initial RoPE $\theta$ across all layers. This approach keeps the base model architecture and other learned weights identical to DRoPE, isolating the effect of the modulation.
By comparing DRoPE’s performance against these shuffled and fixed $\theta$ configurations, we can better attribute performance gains or losses to the effectiveness of the instance-adaptive $\theta_{\text{inst}}$ modulation process.
Appendix F Full Evaluation Results
In this section, we present the detailed results in Section 4.2. Table 12 presents the full results of the zero-shot forecasting experiments conducted at each forecast horizon. For Time-MoE, we report the results as presented in the original paper (Shi et al., 2024).
Table 12: Full results of zero-shot forecasting experiments.
@@ Appendix F Full Evaluation Results
Appendix G Calculation of information density
In this section, we present the formula for the information density illustrated in Figure 1(b). To effectively characterize and compare different time series datasets, we introduce a method to quantify their information density and the variation of this density over time. For this, similar to TimeMixer (Wang et al., 2024a), we utilize spectral entropy, a metric that measures the complexity and compressibility of a signal. In this context, a signal with low spectral entropy, such as one with a few dominant periodic components, is considered to have low information density due to its redundant and predictable nature. Conversely, a signal with high spectral entropy, resembling white noise, has its power spread broadly across the frequency spectrum, indicating a high degree of randomness and thus a higher information density.
Given the observations $\mathbf{x}_{1:T}=(x_{1},\ldots,x_{T})\in\mathbb{R}^{T}$ of a specific dataset, we analyze the time series using a sliding window approach. The series is segmented into $N$ windows, $\mathbf{w}_{i}$, each of size $M$ with a step size of $S$. In our analysis, we use a non-overlapping configuration where both the window size and the step size are set to 128 (i.e., $M=128$, $S=128$). The total number of windows is $N=\lfloor(T-M)/S\rfloor+1$. To mitigate spectral leakage from windowing, we apply a Hamming window to each segment $\mathbf{w}_{i}$:
@@ Appendix G Calculation of information density
Appendix H Showcases
H.1 Showcases of Synthetic Data
Figure 7 showcases representative examples generated by the algorithm using the synthetic dataset. As detailed in Appendix E.1, the generated synthetic dataset is classified into two categories. The first is a composite type, formed from a combination of seasonal, trend, and noise components, which is designated Custom in the figures. The second consists of idealized industrial signals, designated Perfect periodic.

### v3

@@ I.3 Evaluation Datasets
For the evaluation on the TSLib benchmark, the datasets were split into training, validation, and test sets. The split for the ETT and Weather datasets follows the configuration adopted by iTransformer [26]. All other datasets use a 70%/10%/20% ratio for the training, validation, and test sets, respectively.
I.4 Evaluation Length Selection
In this section, we provide further details regarding the selection of historical sequence lengths for the various models evaluated in the TSLib benchmark, supplementing the discussion in Section 4.1.
To accommodate diverse application scenarios, an increasing number of TSFMs [28, 11, 27, 40] have devoted attention to predicting over long contexts. Consequently, we evaluate Kairos and other TSFMs under a long‑context setting. Specifically, we adopt a context length of 2048 time steps and examine four prediction horizons, which are {96,192,336,720}. For TSFMs incapable of processing this context length, we instead employ the context length at which each model achieves its best performance.
For the full-shot deep learning baselines, a hyperparameter search was conducted independently for each dataset to ensure a fair evaluation. We compared the performance of each model using the context length from its original publication against a length of 2048, selecting the superior of the two. To be specific, lengths of 96 and 2048 were compared for DLinear [42], iTransformer [26], TimesNet [41], and Pathformer, while lengths of 336, 512, and 2048 were compared for PatchTST [29].
@@ I.4 Evaluation Length Selection
I.5 Details of DRoPE Analysis
In this section, we explain in more detail the setting of the DRoPE analysis experiment discussed in Section 4.4.2. To more definitively ascertain whether these instance-specific modulations truly capture and leverage beneficial positional information derived from an instance’s FFT features, we designed a series of “shuffle” experiments. These experiments function as a form of causal intervention analysis. The underlying hypothesis is: if the instance-derived $\theta_{\text{inst}}$ modulations are crucial for DRoPE’s performance, then disrupting this linkage by applying $\theta_{\text{inst}}$ modulations from one instance to another (shuffling) should lead to a noticeable degradation in forecasting accuracy.
@@ I.5 Details of DRoPE Analysis
  • Fixed RoPE: To evaluate the architectural contribution of our module, this setting serves as a structural baseline trained from scratch. We replace the DRoPE module with standard RoPE, entirely removing the parameter-predicting MLP. All instances use the static, predefined RoPE $\theta$ across all layers during both training and inference.
By comparing DRoPE’s performance against the inference-time causal interventions (shuffled configurations) and the fixed $\theta$ baseline, we can definitively attribute the performance gains to the effectiveness of the instance-adaptive $\theta_{\text{inst}}$ modulation process.
Appendix J Full Evaluation Results
In this section, we present the detailed results in Section 4.2. Table 16 presents the full results of the zero-shot forecasting experiments conducted at each forecast horizon. For Time-MoE, we report the results as presented in the original paper [34].
Table 16: Full results of zero-shot forecasting experiments. Best results are highlighted in bold, and second best results are underlined. For fair comparison, the full-shot model is excluded from the ranking and is presented for reference only. The dash (-) indicates that the dataset was utilized by the model during its pretraining phase.
@@
Appendix K Calculation of information density
In this section, we present the formula for the information density illustrated in Figure 1(b). To effectively characterize and compare different time series datasets, we introduce a method to quantify their information density and the variation of this density over time. For this, similar to TimeMixer [37], we utilize spectral entropy, a metric that measures the complexity and compressibility of a signal. In this context, a signal with low spectral entropy, such as one with a few dominant periodic components, is considered to have low information density due to its redundant and predictable nature. Conversely, a signal with high spectral entropy, resembling white noise, has its power spread broadly across the frequency spectrum, indicating a high degree of randomness and thus a higher information density.
Given the observations $\mathbf{x}_{1:T}=(x_{1},\ldots,x_{T})\in\mathbb{R}^{T}$ of a specific dataset, we analyze the time series using a sliding window approach. The series is segmented into $N$ windows, $\mathbf{w}_{i}$, each of size $M$ with a step size of $S$. In our analysis, we use a non-overlapping configuration where both the window size and the step size are set to 128 (i.e., $M=128$, $S=128$). The total number of windows is $N=\lfloor(T-M)/S\rfloor+1$. To mitigate spectral leakage from windowing, we apply a Hann window to each segment $\mathbf{w}_{i}$:
@@
Appendix L Showcases
L.1 Showcases of Synthetic Data
Figure 7 showcases representative examples generated by the algorithm using the synthetic dataset. As detailed in Appendix I.1, the generated synthetic dataset is classified into two categories. The first is a composite type, formed from a combination of seasonal, trend, and noise components, which is designated Custom in the figures. The second consists of idealized industrial signals, designated Perfect periodic.